OpenStudy (anonymous):

What are the roots of the polynomial P(s)=3s^2+19 ?

OpenStudy (anonymous):

It does not have a real root

OpenStudy (anonymous):

But it does have roots...

myininaya (myininaya):

3s^2=-19 s^2=-19/3 s=+ or - isqrt{19/3}

OpenStudy (anonymous):

They are called complex

OpenStudy (anonymous):

\[-\sqrt{\frac{19}{3}}i,\sqrt{\frac{19}{3}}i\]

OpenStudy (anonymous):

Thank you to all !!!!!